### v4

@@ Abstract
This study extracts latent factors from the cross-sectional quantiles of firm-level idiosyncratic returns and demonstrates that they carry information that is missed by conventional volatility measures. Notably, exposure to the lower-tail common idiosyncratic quantile factor entails a distinctive risk premium that cannot be explained by existing volatility, downside or tail-related risk factors or characteristics. Furthermore, we demonstrate that factor structures derived from quantiles across the return distribution–which capture its asymmetric features–also possess predictive capabilities regarding aggregate market returns.

### v5

@@ Abstract
We investigate whether the tails of firm-level idiosyncratic return distributions are driven by common shocks. We use quantile factor analysis to extract such common idiosyncratic quantile factors with asymmetric pricing effects and we find a significant premium for innovations to the lower-tail factor: high-beta stocks outperform low-beta stocks by around 7-8% per year. This premium remains significant even when controlling for standard factors, idiosyncratic volatility and tail-risk measures. The downside factor strengthens when intermediary capital is weak and market liquidity is low, and it predicts aggregate market excess returns.